9am a plane leaves Chicago heading west. at 11am a second plane leaves Chicago heading west and traveling 50 mph faster than the first plane. the second plane overtakes the first plane in 5 h. what is the rate of the first plane
Answer by stanbon(75887)

9am a plane leaves Chicago heading west.
at 11am a second plane leaves Chicago heading west and traveling 50 mph faster than the first plane.
the second plane overtakes the first plane in 5 h.
what is the rate of the first plane
-------
2nd plane DATA:
time = 5 hr ; rate = r+50 mph ; distance = r*t = 5r+250 miles
----
1st plane DATA:
time = 5+2 = 7 hr; rate = r mph ; distance = 7r miles
----------------
Equation:
distance = distance
5r+250 = 7r
2r = 250
r = 125 mph (rate of the 1st plane)
======================================
Cheers,
Stan H.
============